Look for PUPD from the following Homogeneous Differential Equations: (-3 +2%)* +(2+2 }dy=0
Answer #1

4-3 +2 ) dx +(67/3 +2) dy =0 let you to y=ux Lorem ut are botas de īs ut 2 du da al 3-20-66²-20 - le- 3-26 6642 px, du (60+2)

(6y²- 4xy - 3x² = 0 This is the solution of differenticel equation, Here, A & C are arbitrary constant. c= 1 / 2